4.13 Let N be a geometric random variable with parameter p - 1/3. Calculate Pr[N s 2], PrN-2|, and PrN22
Answer #1

Here,

N \sim Geometric( p = \frac{1}{3})

The probability mass function of N is

P(N=n) = \frac{1}{3}*(1 - \frac{1}{3})^{n-1},n = 1,2,3,....

Now,

P(N\leq 2) = P(N = 1) + P(N = 2) = 0.333333 + 0.222222 = 0.555555

P(N = 2)= \frac{1}{3}*(1 - \frac{1}{3})^{ 2 - 1}= 0.222222

P(N \geq 2) = 1 - P(N < 2) = 1 - P(N = 1) = 1 - 0.333333 = 0.666667